Methane gas, CH4, effuses through a barrier at a rate of 0.568 mL/minute. If an unknown gas effuses through the same barrier at a rate of 0.343 mL/minute, what is the molar mass of the gas?

**Please explain question **this question has already been posted but the answer was not very explicative.

Thank you!
Answer #1

Rate of effusion is inversely proportional to square root of molar mass

rate = k*sqrt(1/M)

M is molar mass

If we take ratio, above expression becomes

rate(CH4)/rate(unknown) = sqrt (M(unknown)/M(CH4))

we have:

M(CH4) = 16.042 g/mol

rate(CH4) = 0.568

rate(unknown) = 0.343

rate(CH4)/rate(unknown) = sqrt (M(unknown)/M(CH4))

0.568/0.343 = sqrt (M(unknown)/16.042)

sqrt (M(unknown)/16.042) = 1.656

(M(unknown)/16.042) = 2.742

M(unknown) = 43.99 g/mol

Answer: 44.0 g/mol